# ReelWave: A Multi-Agent Framework Toward Professional Movie Sound Generation ![[Uncaptioned image]](x1.png)

### v1

@@ 3.3 Audio-conditioning for on-screen sound
We use three time-varying interpretable signals, namely Loudness, Pitch and Timbre, as the control signals for synchronized on-screen sound generation, drawing inspiration from Sketch2Sound [7], where these sound properties have been shown to achieve impressive results on audio style control and transfer. We further extended the application to video-conditioned audio generation. Since Sketch2Sound is not open-sourced yet, we made some adjustments in our implementation compared to the original paper. Below is a brief summary on how we extract these properties:
@@
Pitch: We extract the raw pitch probabilities of using CREPE (“tiny” variant from PyTorch implementation111https://github.com/maxrmorrison/torchcrepe) [15] pitch estimation model. We further zero out all probabilities below 0.1 in the pitch probability matrix, and finally decode back to the pitch signal.
Timbre: We use the center of mass of the frequency spectrum per audio frame (Spectral Centroid) as the measurement of timbre. Audio that has a brighter timbre will contain frames of a higher spectral centroid. Following [7] we extract and convert the spectral centroid to a MIDI-like representation.
Figure 5 illustrates the conditioning mechanism: after extracting the audio properties, we concatenate them resulting in a conditional vector of shape , where denotes the sequence length.
We train a prediction model to predict the control signals from the video features. The prediction model contains two transformer blocks, three linear blocks, and three prediction heads, where the prediction heads share the same input features. Consider the different length between the video features and the control signals, we use interpolation to reshape the ground truth signals and use MSE loss when training the model. During inference, we again use interpolation to obtain the predicted signals.
@@
The cross-attention mechanism is detailed as follows: Given a query feature , text features and control signals , the output for combining two types of cross-attention is defined as:
The text feature is obtained from the key event description using the original text encoder. During inference, we introduce another parameter for balancing the guidance from text and control signal:
@@ 3.4 Implementation Details
During training, we randomly drop the text feature and control signal with equal probability of 0.05, following [37]. We train separately the prediction model and the conditional layers. To stabilize the training process of the conditional layers, we perform median filtering on the input control signals. During inference, we set as default. The median filter is not used during inference; we empirically find that not using it during inference improves the result.
@@
We employ SynchFormer [13] to extract the video features, which is designed and trained to encode motion dynamics and semantics. We freeze the video encoder during training. Since SynchFormer is trained on 4.8-sec clips, our training and evaluation are conducted on 5-sec videos. The generation backbone for Foley Artist is Auffusion [35] (for on-screen sound) and Stable-Audio-1.0 [6] (for off-screen sound). The backbone for Composer and Voice Actor are MusicGen [4] and Bark [28] respectively. For the video understanding model, we employ LLaVA-Video-7B-Qwen2 [41] for its comparably compact size and powerful performance. We use GPT-4 Turbo [1] in our multi-agent system.
4 Experiments
4.1 Datasets
@@
We use AVSync15 [39] for training and evaluation for the on-screen sound generation module. AVSync15 is a curated dataset from VGGSound Sync [2] that has 1500 high video-audio alignment pairs, which is ideal for training and demonstrating temporal alignment between video and audio. We repeat and concatenate all the videos into 5-sec videos and extract the audio ground truth at 16000 Hz.
4.2 Evaluation Metrics
@@
We employ a comprehensive set of evaluation metrics to evaluate the audio quality and the synchronization for on-screen audio generation. We use the Frechet distance (FD) and Mean KL Divergence (MKL) to evaluate audio fidelity. FD evaluates distributional similarity to evaluate the fidelity of the synthesized audio. MKL measures paired sample-level similarity using mean KL-divergence. To quantify audio-video relevance, we utilize CLIP Score and AV-Align. The CLIP Score is obtained using Wav2CLIP as the audio encoder and CLIP as the video encoder, which quantifies the similarity between the input video and the synthesized audio embeddings in the same representation space.
@@
To evaluate synchronization, we employ # Onset Accuracy, Onset AP, Energy MAE and AV-Align. AV-Align is based on detecting energy peaks in audio-visual modalities, The onset signal is obtained by thresholding the amplitude gradient. As noted in [14], these metrics may ignore relatively quiet sound effects or natural sounds. We further follow [14] to report mean absolute error (MAE) of the energy signals and AV-align score, which is based on detecting energy peaks in audio-video pairs.
4.3 Evaluation and comparison
@@
Table 1 shows that for semantic alignment and audio quality, our model outperforms other baselines on MKL, while consistently ranked second on other metrics. This is a strong indication that ReelWave can successfully predict audio attributes inherent in the given video, offering an explainable temporal and semantic meaning. With only training the model on a small dataset (1,500 samples), our model achieves a reasonable result, while in comparison, Seeing and Hearing, Im2wav use VGGSound Dataset (200,000 samples), and Diff-Foley uses AudioSet (390,000 samples) in addition to VGGSound. ReelWave demonstrates effectiveness and potential to be further extended in larger datasets.
@@
For temporal synchronization, ReelWave is ranked first on energy MAE indicating its better structure prediction, as presented in Table 2. Although our model is ranked average in Onset detection, note that all the baselines are trained on VGGSound dataset, which is a superset of AVSync. While Im2wav achieves excellent results on Onset Acc and AV-align, it is ranked last on others. In summary, our work achieves balanced results across all the metrics, with our goal being to produce on- and off-screen sound.
@@
5 Conclusion and Discussion
Limitation and Future Work. ReelWave uses SynchFormer as the video feature encoder, which performs optimally when the input video does not exceed 5 seconds. While some previous methods use CLIP to extract per-frame features to get longer feature sequence, it has been noted in [14] that this method does not suit the temporal synchronization task. It is thus an important direction to conduct research in designing a better video feature extractor tailored for on-screen sound generation. In addition, ReelWave may be extended for lip synchronization and inferring dialogue from the talking-head video, given stronger video understanding model and the text-to-speech model. This may make a powerful impact in video-to-speech generation.
Conclusion. We present ReelWave, a multi-agent framework for multi-scene audio generation. Our key idea lies in discomposing the generation target to on-screen sound and off-screen sound, where the final audio is produced by overlapping soundtracks with appropriate volume and time occurrence to achieve rich content. Our work also proposes a methodology that uses time-varying interpretable control signals that are derived from the inherent audio. Quantitative results show that even without large-scale datasets, the on-screen sound prediction model can achieve a promising result in terms of semantic alignment, audio quality, and temporal synchronization.

### v2

@@ 3.3. Audio control signals prediction from video
We use three time-varying interpretable signals, namely Loudness, Pitch, and Timbre, as the control signals for synchronized on-screen sound generation, drawing inspiration from Sketch2Sound (García et al., 2024), where these sound properties have been shown to achieve impressive results on audio style control and transfer. We further extended the application to video-conditioned audio generation, made adjustments in our implementation compared to the original paper, as Sketch2Sound is not open-sourced at the time of writing. The following summarizes how we extract these properties:
@@
Pitch: We extract the raw pitch probabilities of using CREPE (“tiny” variant from PyTorch implementation111https://github.com/maxrmorrison/torchcrepe) (Kim et al., 2018) pitch estimation model. We further zero out all probabilities below 0.1 in the pitch probability matrix, and finally decode back to the pitch signal.
Timbre: We use the center of mass of the frequency spectrum per audio frame (Spectral Centroid) as the measurement of timbre. Audio that has a brighter timbre will contain frames of a higher spectral centroid. Following (García et al., 2024) we extract and convert the spectral centroid to a MIDI-like representation.
Figure 5 illustrates the conditioning mechanism: After extracting the audio properties, we concatenate them, resulting in a conditional signal of shape , where denotes the sequence length.
We train a prediction model to predict the control signals from the video features. The prediction model contains two transformer blocks, three linear blocks, and three prediction heads, where the prediction heads share the same input features. Considering the different lengths between the video features and the control signals, we use interpolation to reshape the ground truth signals. During inference, we again use interpolation to obtain the predicted signals.
We use MSE loss when training the model. To stabilize the fluctuation of the conditional signal between each temporal window, we apply a median filter on the ground truth signal during training. During inference, no median filter is used, and we empirically find that it increases the performance. As noted by (Jeong et al., 2024), conventional approaches that use CLIP (Radford et al., 2021) for video encoding may result in suboptimal performance on temporal alignment. We employ SynchFormer (Iashin et al., 2024) to extract the video features, which is designed and trained to encode motion dynamics and semantics. We freeze the video encoder during training. The learning objective is:
where denotes the prediction model that we train, denotes the extracted video feature from SynchFormer, and denotes the conditional signal.
3.4. Signal-conditioning for on-screen sound
@@
The cross-attention mechanism is detailed as follows: Given a query feature , text features , and control signals , the output for combining two types of cross-attention is defined as:
The text feature is obtained from the key event description using the original text encoder. In short, we only add two parameters and on top of each text cross-attention layer, and we use the same query for control signal cross-attention as for text cross-attention. During inference, we introduce another parameter for balancing the guidance from text and control signal:
@@ 3.5. Implementation Details
During training, we randomly drop the text feature and control signal with equal probability of 0.05, following (Ye et al., 2023). We train separately the prediction model and the conditional layers. During inference, we set as default.
@@
Since the video encoder that we use, SynchFormer, is trained on 4.8-sec clips, our training and evaluation are conducted on 5-sec videos. In our multi-agentic, multimodal LLM conversation framewkork, the generation backbone for Foley Artist is Auffusion (Xue et al., 2024) for on-screen sound, and Stable-Audio-1.0 (Evans et al., 2024) for off-screen sound. On the other hand, the backbone for Voice Actor is LipVoicer (Yemini et al., 2023) (for on-screen sound) and Bark (Suno, 2023) (for off-screen sound). The backbone for Composer is MusicGen (Copet et al., 2023). For the video understanding model, we employ LLaVA-Video-7B-Qwen2 (Zhang et al., 2024c) for its comparably compact size and powerful performance. We use GPT-4 Turbo (Achiam et al., 2023) in our multi-agentic system.
4. Experiments
4.1. Datasets
@@
We use AVSync15 (Zhang et al., 2024b) for training and evaluation for the on-screen Foley sound generation module. AVSync15 is a curated dataset from VGGSound Sync (Chen et al., 2021) that has 1500 high video-audio alignment pairs, which is ideal for training and demonstrating temporal alignment between video and audio. We repeat and concatenate all the videos into 5-sec videos and extract the audio ground truth at 16000 Hz.
4.2. Evaluation Metrics
@@
We employ a comprehensive set of evaluation metrics to evaluate the audio quality and the synchronization for on-screen Foley audio generation. We use the Frechet distance (FD) and Mean KL Divergence (MKL) to evaluate audio fidelity. FD evaluates distributional similarity to evaluate the fidelity of the synthesized audio. MKL measures the paired sample-level similarity using mean KL-divergence. To quantify audio-video relevance, we utilize CLIP Score and AV-Align. The CLIP Score is obtained using Wav2CLIP as the audio encoder and CLIP as the video encoder, which quantifies the similarity between the input video and the synthesized audio embeddings in the same representation space.
@@
To evaluate synchronization, we employ # Onset Accuracy, Onset AP, Energy MAE, and AV-Align. AV-Align is based on detecting energy peaks in audio-visual modalities, The onset signal is obtained by thresholding the amplitude gradient. As noted in (Jeong et al., 2024), these metrics may ignore relatively quiet sound effects or natural sounds. We further follow (Jeong et al., 2024) to report the mean absolute error (MAE) of the energy signals and AV-align score, which is based on detecting energy peaks in audio-video pairs.
4.3. Evaluation and comparison
@@
Table 1 shows that for semantic alignment and audio quality, our model outperforms other baselines on MKL, while consistently ranking second on other metrics. This is a strong indication that ReelWave can successfully predict audio attributes inherent in the given video, offering an explainable temporal and semantic meaning. With only training the model on a small dataset (1,500 samples), our model achieves a reasonable result, while in comparison, Seeing and Hearing, Im2wav use VGGSound Dataset (200,000 samples), and Diff-Foley uses AudioSet (390,000 samples) in addition to VGGSound. ReelWave demonstrates effectiveness and potential to be further extended in larger datasets.
@@
For temporal synchronization, ReelWave is ranked first on energy MAE, indicating its better structure prediction, as presented in Table 2. Although our model is ranked average in Onset detection, note that all the baselines are trained on the VGGSound dataset, which is a superset of AVSync. While Im2wav achieves excellent results on Onset Acc and AV-align, it is ranked last on the others. In summary, our work achieves balanced results across all the metrics, with our goal being to produce on- and off-screen sound.
@@
5. Conclusion and Discussion
Limitation and Future Work. ReelWave uses SynchFormer as the video feature encoder, which performs optimally when the input video does not exceed 5 seconds. While some previous methods use CLIP to extract per-frame features to get a longer feature sequence, it has been noted in (Jeong et al., 2024) that this method does not suit the temporal synchronization task. It is thus an important direction to conduct research in designing a better video feature extractor tailored for on-screen sound generation.
Conclusion. We present ReelWave, a multi-agent framework for multi-scene audio generation. Our key idea lies in decomposing the generation target to on-screen sound and off-screen sound, where the final audio is produced by overlapping soundtracks with appropriate volume and time occurrence to achieve rich content. Our work also proposes a methodology that uses time-varying interpretable control signals that are derived from the inherent audio. Quantitative results show that even without large-scale datasets, the Foley on-screen sound prediction model can achieve a promising result in terms of semantic alignment, audio quality, and temporal synchronization.